# VisRet: Visualization Improves Knowledge-Intensive Text-to-Image Retrieval

### v2

@@ 5.2 VisRet Improves T2I Retrieval
Compared with VisRet, the text-to-text retrieval baselines exhibit inconsistent performance across the benchmarks. Captioning the entire corpus with BLIP only works on COCO, which has the closest relationship to its training data, while it does not improve performance over cross-modal retrieval on the other benchmarks with more knowledge intensive queries. We hypothesize that this is caused by the high information density in images, which incurs an inevitable loss of potentially useful information during captioning. On the other hand, with access to cross-modal retrieval results and by directly operating on them, Captioning Reranking and VISA outperform query-agnostic corpus captioning, yet cannot outperform direct cross-modal retrieval by a large margin. Furthermore, these methods can only process a limited number of top-k retrieved items as their cost grows linearly with k. As a result, their performance is limited by the first round of retrieval, leading to near-zero performance on INQUIRE-Rerank-Hard and COCO-Hard.
Finally, comparing VisRet with its single-image variant, we find that using only one generated image as the query slightly harms the performance but still outperforms most of the other baselines. The finding holds for both the lightweight CLIP embedding or the stronger E5-V embeddings, indicating the flexibility of VisRet. Table˜3 presents several examples to demonstrate how the visualization step successfully captures subtle visual features implied by the original text query. As shown in the examples, for a number of ground truth images, the main entity is either in a particular posture that reveals the required information (Visual-RAG and Visual-RAG-ME) or in a complex relationships between other entities (INQUIRE and COCO), posing a challenge to cross-modal similarity matching to recognize and satisfy these constraints. On the other hand, VisRet infuses these constraints naturally into the image query and significantly reduces the workload of the downstream embedding model.
@@ 5.3 VisRet improves Downstream QA
Retrieval Strategy LLM R@1 R@10 R@30 N@1 N@10 N@30
Original Query - 0.210 0.583 0.737 0.210 0.355 0.385
Llama 3.3 70B Instruct 0.240 0.575 0.742 0.240 0.377 0.399
GPT-4o 0.238 0.586 0.737 0.238 0.360 0.395
Llama 3.3 70B Instruct 0.256 0.627 0.790 0.256 0.413 0.437
GPT-4o 0.251 0.645 0.793 0.251 0.431 0.438
To assess the utility of VisRet in real-world applications, we evaluate its downstream VQA accuracy in a RAG pipeline. We compare three settings: (1) using only the model’s internal knowledge, (2) RAG with original text query-based retrieval and (3) RAG with VisRet. Figure˜2 shows the QA accuracy on Visual-RAG and Visual-RAG-ME using GPT-4o as the LVLM reader and CLIP as the retriever. The original query results in low-quality retrieval augmentation, even slightly harming performance on Visual-RAG in top-1 retrieval setting compared to no retrieval. By contrast, VisRet significantly improves QA accuracy in both top-1 and top-10 settings on both benchmarks, boosting accuracy to 0.538 on Visual-RAG and 0.700 on Visual-RAG-ME. Remarkably, on Visual-RAG-ME, VisRet outperforms top-10 retrieval in the original query setting with only top-1 retrieval, highlighting its high accuracy in retrieving the images that contain the required features. Overall, the results confirm that VisRet not only improves retrieval accuracy but also leads to tangible gains in downstream VQA performance. In Section˜6.3, we further demonstrate that VisRet can bring similar performance gains to other models as the VQA reader. In Section˜6.4, we analyze the performance of directly using the generated images as the context and find that while T2I generation improves retrieval, it still cannot replace the retrieved natural images in most cases.
@@ 6 Analyses
In this section, we provide more comprehensive analyses to investigate the effectiveness of VisRet from more perspectives, including the choices of T2I instruction LLM, T2I generation model, and the downstream VQA reader. Finally, inspired by the generative retrieval literature, we conduct a pilot study of whether the generated images could be directly used as the knowledge context for downstream question answering.
6.1 Query Generation Model Choice
Does VisRet work well with other LLMs as the T2I instruction generator? In Table˜4, we study two differently sized open-weight LLMs for rephrasing the query and generating the T2I instruction: Llama 3.1 8B Instruct and Llama 3.3 70B Instruct (Grattafiori et al., 2024). Overall, we observe promising results: For all three LLMs, using the LLMs themselves to generate T2I instructions for VisRet outperforms using the LLM themselves to rephrase the query. While more expensive LLMs achieve higher performance, the small 8B Llama model can already achieve decent performance at a similar level as GPT-4o.
@@ 6.1 Query Generation Model Choice
Table 5: nDCG@k retrieval results on Visual-RAG-ME for different T2I generation models used in VisRet. We report N@1, N@10, N@30, and their average (Avg) for two retrievers (CLIP and E5-V). Best numbers within each retriever are boldfaced. Gemini stands for the Gemini-2.5-Flash-Image-Preview model.
@@
6.2 T2I Generation Model Choice
How strong does the T2I generation model need to be for VisRet to work well? We compare the default T2I generation model (Image-1 with high quality setting) with several other commonly used strong T2I generation models:
  • Stable Diffusion 3 (Esser et al., 2024).
  • FLUX-1.dev (Black Forest Labs, 2024).
  • Qwen-Image (Wu et al., 2025a).
  • DALL-E 3 API (OpenAI, 2023a).
  • Gemini-2.5-Flash-Image-Preview API (Fortin et al., 2025).
@@ 6.2 T2I Generation Model Choice
Table˜5 shows the results with GPT-4o as the T2I instruction generation model and CLIP as the retriever model. We observe that while all the models can improve over cross-modality retrieval to some extent in the VisRet pipeline, their effectiveness increases as the image quality retrieves. The best performance is achieved by the latest proprietary models Image-1 and Gemini-2.5-Flash-Image-Preview. Together, these results suggest that a good T2I generation model with strong instruction following ability is necessary for VisRet. As further T2I generation methods improve, we anticipate that building more cost-efficient versions of VisRet is viable and promising.
6.3 Downstream VQA Model Choice
@@
While we have shown the benefit of VisRet on VQA for GPT-4o, does the improvement hold across LVLMs with different capabilities? In Table˜6, we repeat the VQA experiments with two additional LVLMs: GPT-4o-mini (version gpt-4o-mini-2024-07-18) and GPT-4.1 (version gpt-4.1-2025-04-14). Overall, we observe similar trends as those presented in Figure˜2. Both direct T2I retrieval and VisRet outperform only relying on the model’s knowledge, with VisRet substantially outperforming the former. These results form the foundation for VisRet as a general plug-and-play method to enhance RAG pipelines that rely on accurate T2I retrieval.
6.4 Image Queries as Knowledge
@@
As demonstrated by previous results, a T2I generation model with strong ability to follow instructions and generate realistic images is crucial to the success of VisRet. A natural question is whether it is still necessary to perform retrieval instead of directly using the generated images as the knowledge? In Table˜7, we compare the performance of using a single image as the context with VisRet. Overall, we observe mixed results. For Visual-RAG, GPT-4o-mini achieves slightly higher performance with the generated image than top-1 retrieval, but GPT-4o and GPT-4.1 exhibit the reverse pattern. For Visual-RAG-ME, both GPT-4o-mini and GPT-4.1 prefer the generated image over top-1 retrieval (and even top-10 retrieval). However, when provided with top-10 retrieved images, the models generally exhibit a higher VQA performance than when using the generated image. Therefore, we conclude that retrieving natural images is still crucial for challenging VQA tasks like Visual-RAG and Visual-RAG-ME and cannot be fully replaced by pure T2I generation at this stage. An important further direction is to combine image generation and image retrieval to improve the quality of the retrieved knowledge.
7 Conclusion
Ethics Statement
Potential Risks
Artifact Release
AI Assistant Use
Acknowledgment
Supplementary Material: Appendices
Appendix A Benchmark Data Details
In this section, we present the details of data processing for Visual-RAG, INQUIRE-Rerenk-Hard, and COCO-Hard.
A.1 Visual-RAG
Visual-RAG (Wu et al., 2025b) releases 391 queries with associated image names from iNaturalist 2021 (Horn et al., 2021) and corresponding retrieval labels [^2] https://github.com/visual-rag/visual-rag. To prepare the data, we download the original iNaturalist 2021 dataset and re-collect the images from the train and test set. We were able to identify all annotated images in Visual-RAG except a single image due to a likely path error.
A.2 INQUIRE-Rerank-Hard
To prepare INQUIRE-Rerank-Hard, we accessed the publicly released INQUIRE-Rerank (Vendrow et al., 2024) benchmark [^3] https://huggingface.co/datasets/evendrow/INQUIRE-Rerank. The original test set contained 160 queries, each paired with 100 images retrieved by CLIP. In a pilot study, we tested the retrieval performance of off-the-shelf CLIP and E5-V models. Results showed that CLIP can achieve 0.438 Recall@1 while E5-V achieved 0.506 Recall@1. After manually inspecting the data, we found that for a lot of instances, the negative images are not challenging enough and it is often very straightforward to identify the ground truth images. To highlight the challenging questions, we therefore filtered out the questions on which either CLIP or E5-V can achieve a perfect Recall@1. Overall, we observe that the remaining 59 questions require more nuanced image context understanding and a higher level of knowledge of the organisms themselves, with more challenging confounder negative images.
A.3 COCO-Hard
We derive COCO-Hard from the publicly released validation set of Microsoft COCO 2017 (Chen et al., 2015). The original validation set contained 5000 queries, and we directly use the ground truth images themselves to construct the image corpus. In a pilot study, we found that CLIP is able to achieve 0.424 Recall@1 and 0.756 Recall@10, while E5-V has 0.594 Recall@1 and 0.892 Recall@10. A manual analysis shows that the queries of most of the examples in the original COCO dataset are short phrases centered on a single entity. To increase the task’s difficulty, we rank the examples in reverse order by nDCG@30 and retained the hardest 500 instances. We find these hard queries often describe complex spatial relationships among multiple entities.
Appendix B Implementation Details
VisRet: T2I Generation
To project the text query into the image space, we first instruct an LLM to analyze the query and highlight the key visual features in it. The prompts for three benchmarking datasets are shown in Figure˜3, Figure˜4, and Figure˜5, respectively. Then, we wrap the rephrased query in a prompt template ‘‘Generate a small image of the {rephrased_query}’’ to obtain the final instruction for T2I generation. We use the model gpt-4o-2024-08-06 via OpenAI API with temperature = 0 for instruction generation and the gpt-image-1 model for T2I generation with the quality flag set to "high". For generating multiple images for each query, we find that calling the gpt-image-1 API to return multiple images given a single instruction already results in images generated with a high level of diversity. Therefore, we followed this setting in this paper and save further perturbing the instruction as future work.
VisRet: Retrieval
After obtaining the generated visualizations, we encode both the visualized images and the image corpus via an off-the-shelf CLIP [^4] https://huggingface.co/openai/clip-vit-large-patch14-336 or E5-V [^5] https://huggingface.co/royokong/e5-v encoder and perform a similarity search. Cosine similarity is used as the similarity metric. For RRF, we used $\lambda=1$ to merge the rankings from multiple queries. All the retrieval experiments were performed on a local server with Nvidia A100 GPU.
VQA Answer Generation
We slightly modify the prompt in Visual-RAG to use chain-of-thought prompting (Wei et al., 2022). Concretely, the model is asked to always extract visual information, perform reasoning, and conclude its reasoning with self-verification. We show the detailed prompt in Figure˜6 and Figure˜7.
VQA Evaluation
For Visual-RAG and Visual-RAG-ME, we use the same evaluation prompt released by the authors of Visual-RAG (Wu et al., 2025b), as shown in Figure˜8. Since this prompt is already human engineered for evaluating more complex references and long-form answers and the answers of Visual-RAG-ME are short and easy to evaluate, we did not perform additional prompt engineering. We use the same prompt and gpt-4o-2024-08-06 as the LLM judge for all the VQA experiments in this paper.
Baselines
For the LLM rephrase baseline, we use the same prompt for VisRet T2I instruction to highlight the most important feature that the query is seeking. At the early stage of the project, we performed manual tuning on the prompt and found that the best-performing rephrase also serves as the best-performing T2I generation instruction. Therefore, we report the results with the same prompt for the final version for the paper. For the Corpus Captioning baseline, we use the Salesforce/blip-image-captioning-large model accessed at huggingface. We follow the recommended parameters to run inference on local GPUs, except setting the minimum output length to 10 tokens to encourage a more detailed caption. For the VISA Reranking baseline, we followed the original paper’s public implementation [^6] https://github.com/XLearning-SCU/2025-ICML-VISA, using the sample prompt except changing the model to Qwen2.5-VL-72B-Instruct which we empirically find better performance. For the Captioning Reranking baseline, we simplify the VISA pipeline by directly prompting for a caption for each of the top retrieved images using the following prompt:
"Write a caption that describes the appearance and posture of the main subject. The caption should be concise but informative so that it can be used to retrieve the images from images of the same subject in different postures, states, or environments."

### v3

@@ 5.2 VisRet Improves T2I Retrieval
Text-to-text baselines behave less consistently across benchmarks. Corpus captioning with BLIP helps on COCO, which is closest to its training distribution, but does not improve over cross-modal retrieval on benchmarks with more knowledge-intensive queries. We hypothesize that this reflects the high information density in images and the corresponding loss of potentially useful details during captioning. In contrast, reranking approaches that operate on cross-modal retrieval results, including Captioning Reranking and VISA, outperform query-agnostic corpus captioning. However, they still do not outperform direct cross-modal retrieval by a large margin. These approaches are also constrained by cost, since they process only top-$k$ candidates and their cost grows linearly with $k$. As a result, their performance is limited by the initial retrieval stage, leading to near-zero performance on INQUIRE-Rerank-Hard and COCO-Hard.
@@ 5.3 VisRet Improves Downstream QA
To assess VisRet in a RAG pipeline, we evaluate downstream VQA accuracy under three settings: (1) answering with only the model’s internal knowledge, (2) RAG using retrieval from the original text query, and (3) RAG using VisRet. Figure˜2 reports QA accuracy on Visual-RAG and Visual-RAG-ME with GPT-4o as the LVLM reader and CLIP as the retriever. The original query often provides low-quality retrieval augmentation, and it slightly reduces performance on Visual-RAG in the top-1 retrieval setting relative to no retrieval, suggesting that inaccurate evidence can be more harmful than no evidence.
In contrast, VisRet improves QA accuracy in both top-1 and top-10 settings on both benchmarks, raising accuracy to 0.538 on Visual-RAG and 0.700 on Visual-RAG-ME. Notably, on Visual-RAG-ME, VisRet with top-1 retrieval outperforms the original-query setting even with top-10 retrieval, highlighting its ability to retrieve images that contain the required features with high precision. Overall, these results show that improvements in retrieval quality translate into tangible gains in downstream VQA performance.
@@ 6 Analyses
6.1 Robustness across modules
When VisRet is used in a RAG pipeline, four modules affect end-to-end performance: (1) visualization instruction generation, (2) query visualization, (3) retrieval, and (4) answer generation with an LVLM. We conduct analyses to probe the robustness of VisRet under variations of these modules. Overall, we find that changing the model for (1), (3), or (4) does not substantially alter VisRet’s effectiveness, while performance is more sensitive to (2), which depends strongly on the capability of the T2I generation model. We first summarize the findings below:
  1. 1.
    Using both 8B and 70B Llama (grattafiori2024llama) for instruction generation instead of GPT-4o, VisRet still outperforms using the LLM itself to rephrase the query (Section˜C.1).
  2. 2.
    While most T2I models improve over cross-modality retrieval, strong proprietary T2I generation models yield substantially larger gains when used in VisRet (Table˜3, Section˜C.2). We provide further discussions below.
  3. 3.
    VisRet continues to deliver consistent gains when using a state-of-the-art embedding model (Table˜9, Section˜C.5).
  4. 4.
    Repeating the VQA experiments on Visual-RAG and Visual-RAG-ME with a stronger and a weaker reader LVLM, we find that VisRet’s retrieval improvements consistently translate into downstream QA gains (Section˜C.3).
  5. 5.
    Although the visualized queries improve retrieval, our preliminary study suggests they still cannot replace retrieved natural images in most cases (Section˜C.4).
@@ 6.1 Robustness across modules
Table 3: nDCG@k retrieval results on Visual-RAG-ME for different T2I generation models used in VisRet. We report N@1, N@10, N@30, and their average. Best numbers within each retriever are boldfaced. Gemini stands for the Gemini-2.5-Flash-Image-Preview model. The detailed experiment setup is presented in Section˜C.2.
T2I Model Dependency
Despite the retrieval performance can be improved with a range of T2I generation models, VisRet is limited by the fact that it generally benefits more from stronger T2I generation models. As shown in Table˜3, the latest models from Google and OpenAI dominates the performance on Visual-RAG-ME by a large margin. Setting a lower quality for OpenAI Image-1 slightly harms the performance as well. What leads to the performance gap? In Figure˜3, we present three representative cases where Image-1 outperforms Dall-E 3 (openai_dalle3) on Visual-RAG-ME. We find three recurring failure modes:
  • Lack of focus As shown in the top example, Dall-E 3 fails to zoom in and highlight the key feature wanted, i.e., a flower cluster.
  • Factuality issues As shown in the middle example, Dall-E 3 fails to confidently render the bulb of the organism queries.
  • Weak instruction following As shown in the bottom example, Dall-E 3 fails to comply with the instruction to show the abdomen of the organism, especially in the first visualization.
Among them, factuality reflects a core capability gap, while the other two issues can often be partially mitigated through prompting. For fairness, we use the same prompt across models in Table˜3, but we expect that practical deployments can further benefit from model-specific prompt tuning.
Figure 3: Examples of T2I generations from Image-1 and Dall-E 3 on Visual-RAG-ME. For these examples, VisRet with Image-1 achieves a nearly perfect retrieval score while VisRet with Dall-E 3 scores close to zero.
[…]
6.2 Effectiveness of Repeated Runs
Is multi-image aggregation necessary for VisRet? As shown in Figure˜4, for most examples, VisRet’s N@30 outperforms at least two of its individual samples, highlighting the benefit of diversifying the query through multiple visualizations. At the same time, Table˜4 shows that using only one sample only slightly degrades overall performance. This is because in many cases, such as example 2 and 3 in Figure˜3, independently generated images share substantial information overlap. In such cases, a single visualization can be an efficient option.
Figure 4: Distribution of counts where three-image VisRet outperforming its individual samples on N@30.
Benchmark # Img N@1 N@10 N@30 Avg
Retriever = CLIP
Visual-RAG 3 0.251 0.431 0.438 0.373
1 0.243 0.416 0.425 0.361
Visual-RAG-ME 3 0.460 0.632 0.605 0.566
1 0.449 0.623 0.602 0.558
INQUIRE-Rerank-Hard 3 0.170 0.452 0.455 0.359
1 0.166 0.432 0.457 0.352
COCO-Hard 3 0.034 0.072 0.108 0.071
1 0.033 0.068 0.105 0.069
Retriever = E5-V
Visual-RAG 3 0.299 0.452 0.461 0.404
1 0.292 0.438 0.451 0.394
Visual-RAG-ME 3 0.560 0.643 0.622 0.608
1 0.525 0.638 0.617 0.593
INQUIRE-Rerank-Hard 3 0.220 0.377 0.425 0.341
1 0.205 0.292 0.426 0.308
COCO-Hard 3 0.042 0.173 0.205 0.140
1 0.039 0.167 0.203 0.136
Table 4: Retrieval results (N@k) across benchmarks and image settings for two retrievers (CLIP and E5-V). Avg is the mean of N@1, N@10, and N@30.
Figure 5: Examples of VisRet failing to outperform the LLM query rewriting baseline. Error patterns span T2I instruction generation, visualization, and retrieval failure.
6.3 Success and Failure Examples
We conclude the paper with a qualitative study. Table˜5 presents examples where the visualization step captures subtle visual requirements implied by the original query. Additional examples are provided in Appendix˜D. In many ground-truth images, the target entity appears in a specific posture that reveals the relevant evidence (Visual-RAG and Visual-RAG-ME), or it participates in complex relationships with other entities (INQUIRE and COCO). These constraints are challenging to be preserved in the embedding and surfaced during the cross-modal similarity matching. By contrast, VisRet incorporates them easily through query visualization, reducing the burden on the downstream embedding model.
In Figure˜5, we show failure cases where VisRet does not outperform the LLM query rewriting baseline. We observe three salient error patterns:
  • Instruction Generation As shown in the top row, retrieving a natural image that contains the answer requires two features: (1) the larva form and (2) the relevant body part being shown. The generated T2I instruction omits (1) and includes only (2). Addressing this issue likely requires prompt adjustments.
  • Visualization As shown in the second row, the instruction correctly identifies that the organism’s underside must be shown, but the T2I model fails to generate a perspective that reliably highlights the feature.
  • Retrieval Even when instruction generation and visualization succeed, retrieval can still fail when the embedding model is sensitive to variations such as background or object appearance. In the bottom-row example, both the background (table versus human hand and ground) and the object form (alive versus dead) differ, which leads to retrieval failure.
7 Conclusion
Ethics Statement
Potential Risks
Artifact Release
Human Annotation
Two authors, who are graduate students studying Natural Language Processing, are the only annotators involved in Visual-RAG-ME annotation. Both annotators are supported by the research stipend and the annotation work counted into the working hours. Consent was obtained from both annotators before benchmark curation. The entire benchmark creation process was automatically determined exempt by the institution’s IRB policy. The annotators actively discussed whenever they encounter ambiguity during annotation and reached agreements before proceeding. After the benchmark annotation, we performed a round of human auditing to ensure no question may cause privacy or ethics concerns.
AI Assistant Use
Acknowledgment
Supplementary Material: Appendices
Appendix A Benchmark Data Details
In this section, we present the details of data processing for Visual-RAG, INQUIRE-Rerank-Hard, and COCO-Hard.
A.1 Visual-RAG
Visual-RAG (visual-rag) releases 391 queries with associated image names from iNaturalist 2021 (DBLP:conf/cvpr/HornCBWBA21) and corresponding retrieval labels [^2] https://github.com/visual-rag/visual-rag. To prepare the data, we download the original iNaturalist 2021 dataset and re-collect the images from the train and test sets. We were able to identify all annotated images in Visual-RAG except for a single image due to a likely path error.
A.2 INQUIRE-Rerank-Hard
To prepare INQUIRE-Rerank-Hard, we accessed the publicly released INQUIRE-Rerank (DBLP:conf/nips/VendrowPSBJABH24) dataset [^3] https://huggingface.co/datasets/evendrow/INQUIRE-Rerank. The original test set contained 160 queries, each paired with 100 images retrieved by CLIP. In a pilot study, we tested the retrieval performance of off-the-shelf CLIP and E5-V models. Results showed that CLIP can achieve 0.438 Recall@1 while E5-V achieved 0.506 Recall@1. After manually inspecting the data, we found that for many instances, the negative images are not challenging enough and it is often very straightforward to identify the ground-truth images. To highlight the challenging questions, we therefore filtered out the questions for which either CLIP or E5-V achieves a perfect Recall@1. Overall, we observe that the remaining 59 questions require more nuanced image context understanding and a higher level of knowledge of the organisms themselves, with more challenging confounding negative images.
A.3 COCO-Hard
We derive COCO-Hard from the publicly released validation set of Microsoft COCO 2017 (mscoco). The original validation set contained 5000 queries, and we directly use the ground-truth images to construct the image corpus. In a pilot study, we found that CLIP is able to achieve 0.424 Recall@1 and 0.756 Recall@10, while E5-V has 0.594 Recall@1 and 0.892 Recall@10. A manual analysis shows that the queries of most examples in the original COCO dataset are short phrases centered on a single entity. To increase the task’s difficulty, we rank the examples in reverse order by nDCG@30 and retain the hardest 500 instances. We find that these hard queries often describe complex spatial relationships among multiple entities.
Appendix B Implementation Details
VisRet: T2I Generation
To project the text query into the image space, we first instruct an LLM to analyze the query and highlight the key visual features in implies. The prompts for three benchmarking datasets are shown in Figure˜6, Figure˜7, and Figure˜8, respectively. Then, we wrap the rephrased query in a prompt template ‘‘Generate a small image of the {rephrased_query}’’ to obtain the final instruction for T2I generation. We use the model gpt-4o-2024-08-06 via OpenAI API with temperature = 0 for instruction generation and the gpt-image-1 model for T2I generation with the quality flag set to "high". For generating multiple images for each query, we find that calling the gpt-image-1 API to return multiple images given a single instruction already results in images with a high level of diversity. Therefore, we followed this setting in this paper and save further perturbing the instruction as future work.
VisRet: Retrieval
After obtaining the generated visualizations, we encode both the visualized images and the image corpus via an off-the-shelf CLIP [^4] https://huggingface.co/openai/clip-vit-large-patch14-336 or E5-V [^5] https://huggingface.co/royokong/e5-v encoder and perform a similarity search. Cosine similarity is used as the similarity metric. For RRF, we use $\lambda=1$ to merge the rankings from multiple queries. All the retrieval experiments were performed on a local server with NVIDIA A100 GPU.
VQA Answer Generation
We slightly modify the prompt in Visual-RAG to use chain-of-thought prompting (wei2022chain). Concretely, the model is asked to always extract visual information, perform reasoning, and conclude their reasoning with self-verification. We provide the full prompt in Figure˜9 and Figure˜10.
VQA Evaluation
For Visual-RAG and Visual-RAG-ME, we use the same evaluation prompt released by the authors of Visual-RAG (visual-rag), as shown in Figure˜11. Since this prompt is already human-engineered for evaluating more complex references and long-form answers and the answers of Visual-RAG-ME are short and easy to evaluate, we did not perform additional prompt engineering. We use the same prompt and gpt-4o-2024-08-06 as the LLM judge for all the VQA experiments in this paper.
Baselines
For the LLM rephrase baseline, we use the same prompt for VisRet T2I instruction to highlight the most important features that the query is seeking. At an early stage of the project, we performed manual tuning on the prompt and found that the best-performing rephrase also serves as the best-performing T2I generation instruction. Therefore, we report the results using the same prompt for the final version for the paper. For the Corpus Captioning baseline, we use the Salesforce/blip-image-captioning-large model downloaded from huggingface. We follow the recommended parameters to run inference on local GPUs, except setting the minimum output length to 10 tokens to encourage a more detailed caption. For the VISA Reranking baseline, we followed the original paper’s public implementation [^6] https://github.com/XLearning-SCU/2025-ICML-VISA, using the sample prompt except changing the model to Qwen2.5-VL-72B-Instruct which we empirically find performs better. For the Captioning Reranking baseline, we simplify the VISA pipeline by directly prompting for a caption for each of the top retrieved images using the following prompt:
"Write a caption that describes the appearance and posture of the main subject. The caption should be concise but informative so that it can be used to retrieve the images of the same subject in different postures, states, or environments."
Retrieval Strategy LLM R@1 R@10 R@30 N@1 N@10 N@30
Original Query - 0.210 0.583 0.737 0.210 0.355 0.385
Llama 3.3 70B Instruct 0.240 0.575 0.742 0.240 0.377 0.399
GPT-4o 0.238 0.586 0.737 0.238 0.360 0.395
Llama 3.3 70B Instruct 0.256 0.627 0.790 0.256 0.413 0.437
GPT-4o 0.251 0.645 0.793 0.251 0.431 0.438
@@ Appendix B Implementation Details
Appendix C Further Analyses
In this section, we provide more comprehensive analyses to investigate the effectiveness of VisRet from additional perspectives, including the choices of T2I instruction LLM, T2I generation model, and the downstream VQA reader LVLM. Finally, inspired by the generative retrieval literature, we conduct a pilot study of whether the generated images could be directly used as the knowledge context for downstream question answering.
C.1 Query Generation Model Choice
Does VisRet work well with other LLMs as the T2I instruction generator? In Table˜6, we study two differently sized open-weight LLMs for rephrasing the query and generating the T2I instruction: Llama 3.1 8B Instruct and Llama 3.3 70B Instruct (grattafiori2024llama). Overall, we observe promising results: For all three LLMs, using the LLMs themselves to generate T2I instructions for VisRet outperforms using the LLMs themselves to rephrase the query. While more expensive LLMs achieve higher performance, the small 8B Llama model can already achieve decent performance at a similar level as GPT-4o.
Retrieval Method Visual-RAG Visual-RAG-ME
N@1 N@10 N@30 N@1 N@10 N@30
Original Query 0.279 0.416 0.420 0.330 0.546 0.543
LLM Rewriting 0.302 0.414 0.420 0.390 0.529 0.560
Corpus Captioning (BLIP) 0.128 0.260 0.310 0.220 0.375 0.418
VISA Reranking (top-30) 0.288 0.431 0.442 0.380 0.547 0.541
Captioning Reranking (top-30) 0.280 0.465 0.496 0.140 0.250 0.257
VisRet 0.335 0.476 0.480 0.550 0.667 0.642
Table 9: Additional evaluation results for Retriever = RzenEmbed-v2-7B. nDCG (N@k) is reported. VisRet still strongly outperforms most baselines in most settings except captioning reranking for Visual-RAG’s N@30.
C.2 T2I Generation Model Choice
How strong does the T2I generation model need to be for VisRet to work well? We compare the default T2I generation model (Image-1 with high quality setting) with several other commonly used, strong T2I generation models:
  • Stable Diffusion 3 (DBLP:conf/icml/EsserKBEMSLLSBP24).
  • FLUX-1.dev (bfl-announcing-2024).
  • Qwen-Image (wu2025qwenimagetechnicalreport).
  • DALL-E 3 API (openai_dalle3).
  • Gemini-2.5-Flash-Image-Preview API (fortin2025-gemini25-flash-image).
@@ C.2 T2I Generation Model Choice
Table˜3 shows the results with GPT-4o as the T2I instruction generation model and CLIP as the retriever model. We observe that while all the models can improve over cross-modal retrieval to some extent in the VisRet pipeline, their effectiveness increases with image quality. The best performance is achieved by the latest proprietary models Image-1 and Gemini-2.5-Flash-Image-Preview. Together, these results suggest that a good T2I generation model with strong instruction-following ability is necessary for VisRet. As further T2I generation methods improve, we anticipate that building more cost-efficient versions of VisRet is viable and promising.
C.3 Downstream VQA Model Choice
@@
While we have shown the benefit of VisRet for VQA for GPT-4o, does the improvement hold across LVLMs with different capabilities? In Table˜7, we repeat the VQA experiments with two additional LVLMs: GPT-4o-mini (version gpt-4o-mini-2024-07-18) and GPT-4.1 (version gpt-4.1-2025-04-14). Overall, we observe similar trends as those presented in Figure˜2. Both direct T2I retrieval and VisRet outperform relying only on the model’s knowledge, with VisRet substantially outperforming the former. These results form the foundation for VisRet as a general plug-and-play method to enhance RAG pipelines that rely on accurate T2I retrieval.
C.4 Image Queries as Knowledge
@@
As demonstrated by previous results, a T2I generation model with strong ability to follow instructions and generate realistic images is crucial to the success of VisRet. A natural question is whether it is still necessary to perform retrieval rather than directly using the generated images as the knowledge? In Table˜8, we compare the performance of using a single image as the context with VisRet. Overall, we observe mixed results. For Visual-RAG, GPT-4o-mini achieves slightly higher performance with the generated image than top-1 retrieval, but GPT-4o and GPT-4.1 exhibit the reverse pattern. For Visual-RAG-ME, both GPT-4o-mini and GPT-4.1 prefer the generated image to top-1 retrieval (and even top-10 retrieval). However, when provided with top-10 retrieved images, the models generally exhibit higher VQA performance than when using the generated image. Therefore, we conclude that retrieving natural images is still crucial for challenging VQA tasks like Visual-RAG and Visual-RAG-ME and cannot be fully replaced by pure T2I generation at this stage. An important direction for future work is to combine image generation and image retrieval to improve the quality of the retrieved knowledge.
C.5 Stronger Embedding Models
In Table˜9, we provide additional results using RzenEmbed-v2-7B [^7] https://huggingface.co/qihoo360/RzenEmbed, which is one of the best models on MMEB at the time of writing (2025 December). VisRet still strongly outperforms most baselines in most settings except captioning reranking for Visual-RAG’s N@30. This result highlights both weaker and stronger embedding models benefit from query visualization.